What mass of a non-volatile solute urea (NH2CONH2) need to be dissolved in
100g of water in order to decrease the vapour pressure of water by 25%? also
calculate the molality of the solution.

Grade:upto college level

1 Answers

Nikhil
7 years ago
Po– P/Po = w/3 /w/m+W/MHere, w and m are wt. and molecular wt. of solute, W and M are wt. and molecular weight of solventp = pressure of solution;po = Normal vapour pressureLet the initial (normal) pressure (po) = p∴ Pressure of solution = 75/100 * p = 3/4 pM = 60, M = 118, W = 100 gm∴ p-3/4 p/p = w/60 /w/60 + 100/181/4 = w/60 /(w/60) + 5.55Or 4w/60 = w/60 + 5.553w/60 = w/20 = 5.55Or w = 111 gMolality = No. of moles of solute/Wt. of solvent * 1000= 111 *1000/60 *100 = 18.52 m